OpenStudy (anonymous):

use the multiplication property of inequalities to solve the inequality. graph the solution set. -1/9t ≥ 2

OpenStudy (zzr0ck3r):

multiply each side by -9 and flip the inequality because of the negative

OpenStudy (anonymous):

so t ≤ -4.5?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

t ≤ -18

OpenStudy (anonymous):

sorry I divided before

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

Yes, and to graph use a closed point and arrow headed to the left of -18

OpenStudy (zzr0ck3r):

wait is this \[\frac{1}{9t}or\frac{1}{9}t\]?

OpenStudy (anonymous):

\[\-frac{ 1 }{9}t\]

OpenStudy (anonymous):

sorry negative 1 over 9 and then 7

OpenStudy (anonymous):

\[\frac{ 1 }{ 9 }t\]

OpenStudy (anonymous):

does that change anything?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok thanks

OpenStudy (zzr0ck3r):

:)